# Dynamical dark energy in the light of DESI 2024 data

### v2

@@ 3.5 Comparison with $\Lambda CDM$
We use the Bayes factor concept to determine the preference of this model over the $\Lambda CDM$ model. Bayes factor is calculated as $\ln B_{w\Lambda}=\ln\mathcal{Z}{w}-\ln\mathcal{Z}\Lambda$, where $\mathcal{Z}$ represents Bayesian evidence and the suffixes $w$ and $\Lambda$ represent the current $wCDM$ and $\Lambda$CDM models, respectively.
@@
In general, Jeffrey’s scale is used to find the preference of one model over another model. According to this scale, a negative preference will be considered if $\lvert\ln B_{w\Lambda}\rvert<1$, on the other hand positive, moderate, and strong preferences will be considered if $\lvert\ln B_{I\Lambda}\rvert>1$, $\lvert\ln B_{I\Lambda}\rvert>2.5$, and $\lvert\ln B_{I\Lambda}\rvert>5.0$, respectively  [60]. We have used a publicly available code MCEvidence [34] to directly calculate the Bayes factor from the MCMC chains. From Table.1 it can be seen that the evidence for our model is $4.77$ indicating moderate evidence for the current model over the $\Lambda$CDM model. The preference we find here using the Bayes factor concept is very similar to the evidence obtained for the CPL model in [3]. We also calculated $\delta\xi^{2}_{min}$ for our model in comparison with $\Lambda CDM$, obtaining the value of $-6.11$, which indicates a better fit to data for our proposed model.
Parameter $\Lambda$ CDM $w$CDM
$10^{-2}\omega_{b}$ $2.267^{+0.013}_{-0.014}$ $2.244^{+0.017}_{-0.012}$
$\omega_{cdm}$ $0.116^{+0.00095}_{-0.00071}$ $0.119^{+0.00077}_{-0.0013}$
$H_{0}$ $68.73^{+0.32}_{-0.47}$ $69.91_{-1.2}^{+1.1}$
$\Omega_{DE}$ $0.6917_{-0.0062}^{+0.0065}$ $0.7056^{+0.0061}_{-0.0052}$
$\Omega_{m}$ $0.2967^{+0.0055}_{-0.0044}$ $0.2943^{+0.0052}_{-0.0061}$
$w_{a}$ $-1.27^{+0.45}_{-0.37}$
$w_{b}$ $-0.59\pm 0.20$
$w_{de}$1 $-0.784_{-0.089}^{+0.10}$
$\delta{\chi_{min}^{2}}$ 0 $-6.11$
$\lvert\ln B_{w\Lambda}\rvert$ $4.77$
Table 1: The constraint ($68\%$CL) obtained on cosmological parameters from the combining data sets of Pantheon supernova survey with SH0ES prior  [57], DESI2024 data [3] together with the compressed Planck likelihood.
@@ 4 Conclusions
Recent findings from DESI2024 [3] show a preference for the dynamical nature of dark energy, including a potential transition from phantom to quintessence phases in the recent past. In this study, we introduce an alternative parameterization of the dark energy equation of state, motivated by quintom-type scalar field models, yielding results comparable to those reported in DESI2024 [3]. One interesting feature of our proposed parameterization is that, while it differs from the CPL model at low redshift, it reduces to the CPL form at high redshift. This means that it could provide a richer understanding of cosmology at lower redshifts compared to that of CPL. The model uses trigonometric cosine and hyperbolic cosine functions to represent the equation of state (EoS) for quintessence and phantom nature, respectively.
We used recently released BAO data from [3], Pantheon+ supernova data with SH0ES prior, and a compressed Planck likelihood to constrain the cosmological parameters. Our findings also suggest that the current equation of state (EoS) of dark energy is of the quintessence type, with a possible phantom barrier crossing in the recent past, which aligns with the results of DESI2024[3, 21]. The transition from phantom to quintessence occurs at approximately $z\simeq 0.35$. For this model, we also find that the current value of the Hubble parameter is higher than that in both the CPL and $\Lambda$CDM models. The Hubble tension is significantly reduced in our model, weakening to approximately $2.8\sigma$ compared to the Hubble Space Telescope and SH0ES data [51], and to $1.6\sigma$ when compared with standardized TRGB and Type Ia supernova data [56]. To evaluate its performance at the linear perturbation level, we plotted the temperature anisotropies and the matter power spectrum (MPS) and compared them to the $\Lambda$CDM model. A slight deviation from the $\Lambda$CDM model was observed at high multipoles and large scales, indicating that our model remains consistent even at the perturbation level. We also calculated the Bayes factor to assess the preference of this model over the $\Lambda$CDM model and found a moderate preference for our model, indicating the dynamical nature of dark energy over the cosmological constant.
@@ 5 Acknowledgement
Figure 1: A triangular plot illustrates the constraints on cosmological parameters for the proposed model $wCDM$ (in red), with the $\Lambda$CDM model (in blue) and CPL model (in green) for comparison.
Figure 2: The figure shows two-dimensional posterior plots on $(\Omega_{0},H_{0})$ plane. The value of the $H_{0}$ from SH0ES collaboration [47] is shown by the green band. The deep and light green regions represent $1\sigma$ and $3\sigma$ constraints respectively.
Figure 4: This figure shows the posterior probability $Pr(w|z)$ of the dark energy equation of state (EOS). The dark blue area indicates the $1\sigma$ confidence level, while the light blue areas show the $2\sigma$ confidence level.
Figure 5: This figure shows $Pr(H(z)/(1+z))$ versus $z$ for the current model, compared with Sh0ES survey data  [47] and BAO surveys  [9, 67, 17, 28]. The dark blue area represents the $1\sigma$ confidence level, and the light blue areas represent the $2\sigma$ confidence level.

### v3

@@ 4 Comparison with $\Lambda CDM$ and the satus on the Hubble Tension
Bayes factor is calculated as $\ln B_{w\Lambda}=\ln\mathcal{Z}{w}-\ln\mathcal{Z}\Lambda$, where $\mathcal{Z}$ represents Bayesian evidence and the suffixes $w$ and $\Lambda$ represent the current $wCDM$ and $\Lambda$CDM models, respectively.
@@
In general, Jeffrey’s scale is used to find the preference of one model over another model. According to this scale, a negative preference will be considered if $\lvert\ln B_{w\Lambda}\rvert<1$, on the other hand positive, moderate, and strong preferences will be considered if $\lvert\ln B_{I\Lambda}\rvert>1$, $\lvert\ln B_{I\Lambda}\rvert>2.5$, and $\lvert\ln B_{I\Lambda}\rvert>5.0$, respectively  [71]. We have used a publicly available code MCEvidence [72] to directly calculate the Bayes factor from the MCMC chains. From Table 2 it can be seen that the evidence for our model is $5.066$, indicating strong evidence for the proposed $w_{q}$ model over the $\Lambda$CDM model, whereas the CPL model also shows similar strong evidence over $\Lambda$CDM.
We also calculated $\Delta\chi^{2}_{min}$ for the $w_{q}$ and CPL model in comparison with $\Lambda CDM$ which is given in Table 2. The lowest value of $\Delta\chi^{2}_{min}$ is obtained for the $w_{q}$ model which is $-17.5$, indicating a better fit to data for the proposed model.
We have also used the concept of the Akaike Information Criterion (AIC) to compare the models against $\Lambda$CDM. The AIC is defined as follows:
$\Delta\text{AIC}=\chi^{2}_{\text{min},\mathcal{M}}-\chi^{2}_{\text{min},\Lambda\text{CDM}}+2(N_{\mathcal{M}}-N_{\Lambda\text{CDM}})$ (8)
Here the $\mathcal{M}$ and $N_{\Lambda\text{CDM}}$ represent the number of free parameters of the model $\mathcal{M}$ and the $\Lambda$CDM. In general, a negative $\Delta$AIC when compared to the $\Lambda$CDM model indicates the preference for the model. As it can be seen from Table 2 the lowest $\Delta$AIC corresponds to the $w_{q}$ model suggesting preference for the proposed model over $\Lambda$CDM.
To quantify the status of Hubble tension in the proposed model we have used two different approaches. One is the Gaussian tension estimator and another is the $Q_{DMAP}$ ( difference of the maximum a posteriori) estimator. The estimation of tension we obtained using these two approaches for the $w_{q}$ model is given in Table 2 together with $\Lambda$CDM and CPL models for comparison.
The Gaussian estimator proposed in [73], $T_{H0}=\frac{\left|H_{0}-H_{0}^{loc}\right|}{\sqrt{\sigma_{H0}^{2}+\sigma_{\mathrm{loc}}^{2}}}$, where $T_{H0}$ denotes the estimation of the amount of tension, $H_{0}$ and $H_{0}^{loc}$ is the mean value of the Hubble parameter obtained from the current analysis and local universe observations, respectively, while $\sigma^{2}$ represents the corresponding variance of the posteriors. By comparing our results with the result obtained in  [2] $\left(H_{0}=73.04\pm 1.04\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}\right)$, from the Hubble Space Telescope and the SH0ES data, the tension status is $T_{H0}\simeq$ $2.86\sigma$. On the other hand, when comparing it with the reported value of $\left(H_{0}=73.22\pm\right.$ $2.06\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ ) from the standardized TRGB and Type Ia supernova data sets in  [3], it reduces to $T_{H0}\simeq 1.69\sigma$. If we consider the classification of the tension in Table IV in  [73] for the first case, the tension is moderate and weak for the latter case.
Another tensiometer $Q_{DAMP}$ was proposed in [74] to check how the inclusion of the data from the $SH0ES$ measurement can impact the fit for a proposed model $\mathcal{M}$ and has been used in [75] to test the status of the Hubble tension in for a wide range of models. To estimate it one needs to find the best-fit chi-square between the combined data set with $SH0ES$ observation and without it as $\Delta\chi^{2}=\chi^{2}_{\min,\mathcal{D}+\text{SH0ES}}-\chi^{2}_{\min,\mathcal{D}}$ and the tension will be estimated as $T=\sqrt{\Delta\chi^{2}}$.  [^1] One can only estimate the $Q_{DAMP}$ tension for datasets differing by one degree of freedom. Since we have used the Pantheon+SH0ES data for constraining the models to calculate $Q_{DAMP}$ we have used a similar approach to [76]. We have imposed a Gaussian prior on the calibration of the absolute magnitude $M_{B}$ of the supernovae Pantheon-plus sample. This is done to exclusively compute $Q_{DAMP}$ for the models, and we have verified that this substitution does not alter the results. The value of the $Q_{DAMP}$ for the $\Lambda$CDM mode we obtained is very similar to the one obtained in the [76] (Table III) using the same set of data. In Table 2 we have listed $Q_{DAMP}$ for $w_{q}$ CDM model and $\Lambda$ CDM and CPL models. The $Q_{DAMP}$ for the $\Lambda$CDM model is 5.56 $\sigma$ whereas for the $w_{q}$ it is $4.89\sigma$ showing a slight reduction in the tension. Also, the status of the tension using $Q_{DAMP}$ for the CPL model is very similar to the $w_{q}$ model. From Table. 2 it can be seen that for the $w_{q}$ model, the $\chi_{min}^{2}$ is lowest for the combination of all data sets indicating that this model fits the data better than others. It can be seen from Fig. 5 and Fig.4 that this model mimics the $\Lambda$ CDM at higher redshift and differs from it at lower redshift, improving fitting to the data, hence reducing the tension.
@@ 5 Conclusions
Recent findings from DESI2024 [1] show a preference for the dynamical nature of dark energy, including a potential transition from phantom to quintessence phases in the recent past. In this study, we introduce an alternative parameterization of the dark energy equation of state, motivated by quintom-type scalar field models, yielding results comparable to those reported in DESI2024 [1]. One interesting feature of our proposed parameterization is that, while it differs from the CPL model at very low redshift, it reduces to the CPL form at high redshift ($z<1$), whereas at higher redshift this model mimics the cosmological constant. This means that it could provide a richer understanding of cosmology at lower redshifts compared to that of CPL while maintaining consistency with early-time cosmology. The model uses trigonometric cosine and hyperbolic cosine functions to represent the equation of state (EoS) for quintessence and phantom nature, respectively.
We used recently released DESI BAO R1 data from [1], Pantheon+ supernova data with and without Chepids host distance anchor, and a compressed Planck likelihood to constrain the cosmological parameters. Our findings also suggest that the current equation of state (EoS) of dark energy is of the quintessence type, with a possible phantom barrier crossing in the recent past, which aligns with the results of DESI2024[1, 59]. The transition from phantom to quintessence occurs at approximately $z\simeq 0.35$. For this model, we also find that the current value of the Hubble parameter is higher than that in $\Lambda$CDM models and Hubble tension is significantly reduced in our model, weakening to approximately $2.8\sigma$ compared to the Hubble Space Telescope and SH0ES data [2], and to $1.6\sigma$ when compared with standardized TRGB and Type Ia supernova data [3]. To evaluate its performance at the linear perturbation level, we plotted the temperature anisotropies and the matter power spectrum (MPS) and compared them to the $\Lambda$CDM model. A slight deviation from the $\Lambda$CDM model was observed at high multipoles and large scales, indicating that our model remains consistent even at the perturbation level. We also calculated the Bayes factor and $\Delta AIC$ to assess the preference of this model over the $\Lambda$CDM model and found a strong preference for our model, indicating the dynamical nature of dark energy over the cosmological constant.
@@ 6 Acknowledgement
Appendix A Evolution of the $w_{q}(z)$ for different choices of the $w_{a}$ and $w_{b}$ parameters.
Here we discuss the evolution of $w_{q}(z)$ with the redshift $z$ for different choices of the $w_{a}$ and $w_{b}$ parameters. As mentioned in Section 2, this parametrization can mimic the cosmological constant, $w_{q}\simeq-1$, irrespective of the choice of model parameters as $\frac{a}{a_{0}}\rightarrow 0$. However, during late times, it may exhibit richer behavior. In Fig.11, we have plotted $w_{q}(z)$ vs $z$ for four different choices of the $w_{a},w_{b}$ parameters, each displaying different qualitative behavior of the $w_{q}$ evolution at late times. In two cases, $w_{q}$ evolves from $\simeq-1$ and dips into either the quintessence (in red) or phantom (in green) regions without any phantom barrier crossing. In one case, $w_{q}$ evolves into the quintessence region and currently transitions to the phantom region through a phantom barrier crossing (in yellow). The plot in blue shows the opposite behavior, where $w_{q}$ crosses the phantom barrier from the phantom to the quintessence domain. Our choice of priors on the $w_{a},w_{b}$ parameters in this analysis incorporates all these aforementioned cases.
Figure 11: Evolution of $w_{q}(z)$ with the redshift $z$ for different choices of the $w_{a}$ and $w_{b}$ showcasing four different type of scenarios.